OpenStudy (anonymous):

Solve the equation x^3+4^2-25x-100=0

OpenStudy (watchmath):

\(x^2(x+4)-25(x+4)=0\) \((x^2-25)(x+4)=0\) \(x^2=25\Rightarrow x=\pm 5\) or \(x=-4\)